Prioritizing the requirements

Prioritizing is used to limit the scope and deliver the most essential functionality as early as possile. Both plan-based an agile development approaches view prioritization as a fundamental activity, but they differ in their basic strategies. A requirements prioritization strategy determines what requirements are implemented and in what order they will be implimented in to maximize the progress towards a goal. A goal could be “maximize value”, “minimize cost”, or “minimize effort”. The purpose of POM was find an effective prioritization strategy to maximize the valuecost ratio of implimenting the requirements.

Simulation requirements and assumptions A requirement is represented as a structure containing a cost and a value. A base set of requirements is generated by assigning uniform random values for the cost and value of the requirements. After each iteration requirement volatility is handled by updating each requirement’s value with a normal distributed random value and a Poisson number of new requirements.

Simulation results figure center FIND SOME GOOD TITLE FOR THIS tabular— c — c — c — c — HIGH AG2 AG2 HY, AG

figure Low Dynamism figure:POM1-low-dynamism